# A Super Version of a theorem of Fricke-Klein

### v1

@@ 4. Character Varieties
The ideal $I_{\Gamma}$ is generated by the relations induced by $\Gamma$, which tell you that the images of the generators still have to obey the relations of $\Gamma$, because otherwise it would not be a group homomorphism. By the quotient nature of $\mathcal{R}(\Gamma,G)$ it is not obvious if it is an integral domain, which over $\mathbb{C}$ allows you to identify Specm and Spec. This space algebraically classifies all so-called $G$-representations. However we do want to quotient out all conjugacy equivalent representations. Thus we define
@@
(53) $\operatorname{Spec}(\mathcal{R}(\Gamma,G)^{G})$
@@
If we now only consider reductive algebraic groups then it turns out that $\operatorname{Specm}(\mathcal{R}(\Gamma,G)^{G})=\operatorname{Hom}(\Gamma,G)$ is a good categorical quotient and by Nagata’s theorem it is finitely generated. One should remark, that non-reduced points may still exists. Let $\operatorname{Hom}^{i}(\Gamma,G)$ be the subset of irreducible representations. Then we also have that
@@
Recalling the theorem of Fricke and Klein, we can now reinterpret it as a special case of the character variety of the free group and in particular an explicit solution to the question of finding minimal generators of a character variety.
We can easily see, that there are no obstructions to replacing the algebraic group with an algebraic supergroup. The character variety of $\operatorname{OSp}_{\mathbb{R}}(1|2)$ has already been investigated from the analytic point of view for example in [19], where Penner-coordinates were generalized. It has also appeared in [23], where they gave an analogous recursion formula to Mirzakhani for super hyperbolic surfaces with heuristics coming from supergeometry. The statement was proven with algebro-geometric methods in [18].
@@
Seeing how this construction actually only cares about the Hopf algebras, we can also generalize this question to quantum groups. This will definitely work if we only consider group relations, that can be manipulated in a way such that no inverses appear, because if we consider the quantum group via algebra-valued matrices, then the inverse does not necessarily live in the same algebra. A heuristic construction was already given in [1]. We note, that for quantum groups one reasons purely in the dual picture of the Hopf algebra, so one has to think about coinvariants of the conjugation coaction.
5. Main Result
Let us return from the geometric world and consider purely algebraic world again, where we now state our main result. Let $V$ be a $1|2$-dimensional super vector space.
Theorem 5.1.
@@ 5. Main Result
Let us consider the algebra
(54) $\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}.$
By the first fundamental theorem of the $\operatorname{OSp}(V)$ we know that the number of polynomial invariants of $V^{\otimes 4}$ is 10. By the second fundamental theorem of $\operatorname{OSp}(V)$ we know that there is one relation for $\operatorname{dim}(V)=1|2$ and thus 9 independent generators. Because the multilinear invariants of matrices and vectors are the same, the polynomial invariants are also the same. By reductivity we have
@@
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc,$
together with $\operatorname{Ber}(A)=1$. We know that there exists no odd invariants on $\mathbb{C}[\operatorname{End}(V)^{2}]$ and we also know the co-action on $I$ sends the odd generators to odd generators. Thus all the odd relations, do not give us an invariant of $I$. However, the relation $\operatorname{Ber}(A)=1$ is invariant under the co-action, thus it is an invariant vector in $I$. It follows that
Therefore we have
(57) $\begin{split}\#\text{gen. }\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}&=\#\text{gen. }\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)}+\#\text{gen. }I^{\operatorname{OSp}(V)}\\ 9&=x+2\\ x&=7\end{split}$
By an earlier theorem, we know that all multilinear invariants are given by traces, thus all these invariants are traces as well. ∎
This result can be reinterpreted as the fact that the character variety of the free group on 2 letters of $\operatorname{OSp}(V)$ has 7 independent generators. Which is a deviation of the result for $SL(2,\mathbb{C})$.
There have been results, where in the real analytic case one can prove, that this space is actually of the expected real dimension $6|4$. One can easily see, why this fails in the algebraic case. We have seen that all invariants on the space of matrices, are given by a supertrace. We recall that
and thus we cannot determine any odd generators with the help of traces. We can also interpret as an incarnation of the fact, that there is a natural $\mathbb{Z}_{2}$-action acting on the odd coordinates. Explicitly, it is
$=\begin{pmatrix}1&0&-\xi\\ 0&1&0\\ 0&\xi&1\end{pmatrix}$
So this tells us, that if we take the quotient by the full group $\operatorname{OSp}(1|2)$, that the space has to be of purely bosonic dimension. One also sees by this calculation, that taking the full ring of invariants can only give you a non-reduced ordinary scheme, since the odd degrees of freedom are not invariant, as we always have to pair two odd coordinates, that will be flipped simultaneously to become an invariant.
This means what is $\operatorname{OSp}(1|2)$-invariant is the product of both odd guys. One reckons that one should consider the quotient $\operatorname{{POSp}}(1|2)\coloneqq\operatorname{OSp}p(1|2)/\mathbb{Z}_{2}$ instead.
Making this change, one also needs to consider the character variety of $POSp(1|2)$ instead.
Remark 5.2.
One might be disappointed by the structure of this character variety. However, as it turns out there are also other instances of moduli spaces of superobjects, that are of purely even nature like the picard scheme of a SUSY curve [3] or the coarse moduli of a superstack [4].

### v2

@@ 5.1. Character Varieties
The ideal $I_{\Gamma}$ is generated by the relations induced by $\Gamma$, which tell you that the images of the generators must satisfy the relations of $\Gamma$, because otherwise it would not be a group homomorphism. By the quotient nature of $R(\Gamma,G)$ it is not obvious if it is an integral domain, which over $\mathbb{C}$ allows you to identify Specm and Spec. This space algebraically classifies all so-called $G$-representations. However we do want to quotient out all conjugacy equivalent representations. Thus we define
@@
(58) $\operatorname{Spec}(R(\Gamma,G)^{G})$
@@
If one considers reductive algebraic groups then it turns out that $\operatorname{Specm}(R(\Gamma,G)^{G})=\operatorname{Hom}(\Gamma,G)$ is a good categorical quotient and by Nagata’s theorem it is finitely generated. One should remark, that non-reduced points may still exist. Let $\operatorname{Hom}^{i}(\Gamma,G)$ be the subset of irreducible representations then
@@
We can easily see, that there are no obstructions to replacing the algebraic group with an algebraic supergroup.
@@
Seeing how this construction actually only cares about the Hopf algebras, we can also generalize this question to quantum groups. This will definitely work if we only consider group relations, that can be manipulated in a way such that no inverses appear, because if we consider the quantum group via algebra-valued matrices, then the inverse does not necessarily live in the same algebra. A construction was already given in the physics literature [1]. We note that for quantum groups one reasons purely in the dual picture of the Hopf algebra, so one has to think about coinvariants of the conjugation coaction.
Using this viewpoint, we can reinterpret 4.1 as the fact that the character variety of the free group on 2 letters of $\operatorname{OSp}(V)$ has 7 independent generators, representing a deviation of the classical result for $SL(2,\mathbb{C})$.
There have been results [25], where in the real analytic case one can prove that this space is actually of the expected real dimension $6|4$. One can easily see, why this fails in the algebraic case. We have seen that all invariants on the space of matrices, are given by a supertrace. We recall that
and thus we cannot determine any odd generators with the help of traces. It is also a consequence of the fact that there is a natural $\mathbb{Z}_{2}$-action acting on the odd coordinates. Explicitly, it is
$=\begin{pmatrix}1&0&-\xi\\ 0&1&0\\ 0&\xi&1\end{pmatrix}$.
So this tells us, that if we take the quotient by the full group $\operatorname{OSp}(1|2)$, that the space has to be of purely bosonic dimension. One also sees by this calculation that taking the full ring of invariants can only give you a non-reduced ordinary scheme, since the odd degrees of freedom are not invariant, as we always have to pair two odd coordinates that will be flipped simultaneously to become an invariant.
This means what is $\operatorname{OSp}(1|2)$-invariant is the product of both odd guys. From this perspective, one should consider the quotient $\operatorname{{POSp}}(1|2)\coloneqq\operatorname{OSp}(1|2)/\mathbb{Z}_{2}$ and consequently, one also needs to consider the character variety of $\mathrm{POSp}(1|2)$. The character variety of $\operatorname{OSp}_{\mathbb{R}}(1|2)$ has already been investigated from the analytic point of view in [25], where Penner-coordinates were generalized. It has also appeared in [31], where they gave an analogous recursion formula to Mirzakhani [22] for super hyperbolic surfaces with heuristics coming from supergeometry. The statement was proven with algebro-geometric methods in [24].
Remark 5.5.
One might be disappointed by the structure of this character variety. However, as it turns out there are also other instances of moduli spaces of super objects that are of purely even nature like the Picard scheme of a SUSY curve [4] or the coarse moduli of a superstack [6]. This suggests that trying to obtain a coarse moduli in supergeometry is ill-advised and one should work purely stack theoretically so you do not accidentally kill all odd directions.
5.2. Character Stack
In view of (4.1) it is clear that to get a proper handle on the character variety of algebraic supergroups one should really be looking at the character stack of the algebraic supergroup instead. We give a brief account on a gradedness result of the character stack of $\mathrm{OSp}(1|2)$. Let $X$ be a super Riemann surface and consider the representation variety
(60) $\mathrm{Rep}(X)\coloneqq\operatorname{Hom}(\pi_{1}(X),\operatorname{OSp}(1|2))$
and its stack quotient $\mathfrak{X}\coloneqq[\mathrm{Rep}(X)/\operatorname{OSp}(1|2)]$, called the character stack. For brevity, we will only define the notion of prestacks and argue, why for this papers purpose it is enough and we will give a short user’s guide on how one works with stacks in practice and for more details we refer to [5].
Definition 5.6.
A pre-superstack is an element in the category
(61) $\mathrm{PresStk}\coloneqq\mathrm{Fun}(\mathrm{sRings},\mathrm{Gpd}),$
i.e. it is a $(2,1)$-presheaf of groupoids.
Remark 5.7.
To turn a pre-superstack into a superstack, one needs to equip it with a Grothendieck topology and it needs to satisfy the descent condition for the chosen topology (this means the cover needs to satisfy a higher categorical gluing condition, in this case a 2-categorical gluing condition).
Conceptually, the thory of stacks is best approached trhough the functor of points perspective. From this perspective the theory of stacks does not look much different from the theory of schemes. We will state two facts that allow us to take a shortcut, so we do not have to think about actual stacks. Firstly, the category of quasi-coherent sheaves on the prestack $\mathfrak{X}$ is equivalent to the category of quasi-coherent sheaves on its stackification $\mathfrak{X}^{+}$
(62) $\mathrm{QCoh}(\mathfrak{X})\cong\mathrm{QCoh}(\mathfrak{X}^{+}).$
Furthermore, the category of superstacks forms a full subcategory of the category of pre-superstacks, so it admits a fully faithful embedding, that is essentially surjective
(63) $\mathrm{sStk}\subset\mathrm{PresStk}.$
This tells us that we do not need to care about the topologies of the stack, if you want to write down morphisms of stacks, the only thing one needs to do is to write down a natural transformation between the functor of points of the stacks and in a similar vein, you do not need to think about the topology to be able to think about quasi-coherent sheaves. Moreover, by its classical definition a stack is also a category fibered in groupoids, so in particularly it is a category, so one only needs to write down a functor to construct a morphism of stacks that commutes with projections to the base category. With all this out of the way, we turn to our short analysis of the character stack.
We want to investigate whether this space is graded, since it is one of the most basic properties one can check. The question whethera space is graded or not depends on a classical obstruction theory argument, i.e. does a certain cohomology class vanish on the underlying space [34] [18]. Thus in this specific setting, it is enough to think about the prestack/functor of points of the character stack. We recall that for a stack quotient $[X/G]$ of a scheme, we have an equivalence of categories
(64) $\operatorname{QCoh}([X/G])\cong\operatorname{QCoh}^{G}(X).$
Furthermore, for $X\in\operatorname{CRing}^{\operatorname{op}}$ and $\mathcal{F}\in\operatorname{QCoh}(X)$, $H^{i}(X,\mathcal{F})=0\ \forall i>0$. We can also rephrase this as the global section functor being exact.
Proposition 5.8.
Let $k$ be an algebraically closed field of characteristic 0, $G$ a reductive algebraic supergroup, and $X$ an affine $k$-superscheme. Then
(65) $H^{i}_{G}(X,\mathcal{F})=0\quad\forall i>0.$
Remark 5.9.
This property is called being cohomologically affine.
@@ 5.2. Character Stack
The global sections functor $\Gamma$ is exact, because $X$ is affine and $(-)^{G}$ is exact, because $G$ is reductive. $G$-equivariant sheaf cohomology is defined to be the right derived functor of the composition $(-)^{G}\circ\Gamma$, thus by composition of exact functors, the composition is also exact. We obtain
(66) $H^{i}_{G}(X,\mathcal{F})\subset H^{i}(X,\mathcal{F})=0\ \forall i>0.$
Corollary 5.10.
Let $X\in\operatorname{Sch}$ and $G\in\operatorname{Grp(Sch)}$. If $X$ is cohomologically affine and smooth and the fixed point functor $(-)^{G}$ is exact, then the stack quotient $[X/G]$ is graded.
Proof.
Let $\mathcal{J}$ be the sheaf of nilpotents on $[X/G]$ and $\mathcal{E}\coloneqq\mathcal{J}/\mathcal{J}^{2}$. Furthermore, we denote $\mathcal{F}^{i}\coloneqq T_{(-)^{i}}[X/G]_{\operatorname{red}}\otimes\wedge^{i}\mathcal{E}^{\vee}$. To prove $[X/G]$ is graded, we need to show the that obstruction classes
(67) $\omega_{i}\in H^{1}([X/G]_{\operatorname{red}},\mathcal{F}^{i})$
vanish [34]. Given the canonical inclusion $\iota:[X/G]_{\operatorname{red}}\rightarrow[X/G]$ one can pushforward the sheaves along the inclusion and one can make the following identifications
(68) $H^{1}([X/G]_{\operatorname{red}},\mathcal{F}^{i})=H^{1}([X/G],\iota_{*}\mathcal{F}^{i})=H^{1}_{G}(X,\iota_{*}\mathcal{F}^{i})=0,$
where the last identification comes from the previous theorem. All obstruction classes have to vanish, therefore $[X/G]$ is graded. ∎
As a consequence of gradedness, the superstack can also be written in the following form:
(69) $[X/G]\cong([X/G]_{\operatorname{red}},\wedge^{\bullet}\mathcal{E}),$
where $\mathcal{E}\coloneqq\mathcal{J}/\mathcal{J}^{2}$. As a consequence, we get an easier description of the Berezinian bundle
(70) $\operatorname{Ber}(\mathcal{T}_{[X/G]_{sm}})=\operatorname{det}(\mathcal{T}_{[X/G]_{sm}})\otimes\operatorname{det}^{-1}(\wedge^{top}\mathcal{E})$
Corollary 5.11.
The smooth locus character stack $\mathfrak{X}_{sm}$ is graded.
Given a point $x:\mathrm{Spec}(\mathbb{C})\rightarrow\mathfrak{X}$, we can consider the tangent complex at the point $x$, which is given by the two-term complex
(71) $[\mathfrak{g}\longrightarrow\mathcal{T}_{X,x}].$
Along the smooth locus, we obtain the tangent space
(72) $\mathcal{T}_{X,x}/\mathfrak{g}.$
This object will have the expected dimension of $(2g-2)\mathrm{dim}(\mathrm{OSp}(1|2))$ and for the details on the explicit calculations we will refer the interested reader to A.2.
Appendix A Computations
This appendix provides a detailed derivation of the dimension of the tangent space of the representation variety $\mathcal{T}_{X,x}$. We conclude that the character stack has the expected dimension for a super Riemann surface of genus $g$. The classical proofs can be adapted with minimal modification from the literature, which we will be pointing towards for each statement. The necessary modification one has to do to adapt these results to the supergeometric case is replace the usual dual numbers, by the so-called super dual numbers, which refers to the superalgebra $k[\varepsilon,\eta]/(\varepsilon^{2},\varepsilon\eta)$, where $\mathrm{deg(\varepsilon)=0}$ and $\mathrm{deg}(\eta)=1$. We begin by proving the following lemma.
Lemma A.1 ([21]).
Let $(A,\Delta)$ be a super Hopf algebra over k and let $I$ denote the kernel of the co-identity map $\epsilon$, then $A\cong k\oplus I$ as k-modules.
Proof.
The co-identity map is a morphism of super algebras, so in particular it is a morphism of super vector spaces, hence we get the following sequence
@@ Appendix A Computations
This is a sequence of super vector spacecs, so it splits and we get the wanted isomophism. ∎
The next theorem was originally proven in [17]. However, we will follow the more detailed presentation of [27].
Theorem A.2.
Let $R(\Gamma,G)$ be the coordinate ring of the representation variety $\mathrm{Hom}(\Gamma,G)$ for $\Gamma$ a finitely generated group and $G$ an algebraic supergroup. Then the Zariski tangent space at a point $\rho$ is isomorphic to $Z(\Gamma,Ad_{\rho})$.
Proof.
We can now describe a morphism from the supergroup $G(\mathbb{C}/[\varepsilon,\eta])$ to the Lie algebra $\mathfrak{g}$. Define
$\sigma(g)=g\pi(g)^{-1}$
where $\pi$ is a morphism of supergroups $G(\mathbb{C}[\varepsilon,\eta](\varepsilon^{2},\eta\varepsilon))\rightarrow G(\mathbb{C})$ induced by $\mathbb{C}[\varepsilon,\eta](\varepsilon^{2},\eta\varepsilon)\rightarrow\mathbb{C}$, where we send both variables to zero. By applying $\pi$ again, we can check if it is an element of the Lie algebra:
$\pi(\sigma(g))$ $=\pi((g\pi(g)^{-1}))$
$=\pi(g)\pi^{2}(g^{-1})$
$=e$
We can now compute the following
$\sigma(g_{1}g_{2})$ $=g_{1}g_{2}\pi(g_{1}g_{2})^{-1}$
$=g_{1}g_{2}\pi(g_{2})^{-1}\pi(g_{1})^{-1}$
$=g_{1}(\pi(g_{1})^{-1}\cdot\pi(g_{1}))\sigma(g_{2})\pi(g_{1})^{-1}$
$=\sigma(g_{1})\cdot Ad_{\pi(g_{1})}\sigma(g_{2})$
And we are allowed to switch to additive notation and we get the 1-cocycle condition for group cohomology
$\sigma(gh)=\sigma(g)+Ad_{\pi(g)}\sigma(h)$
So it is a cycle in $Z^{1}(\Gamma,\mathfrak{g})$.
We now also get a map from $\Gamma\rightarrow\mathfrak{g}$, by using the universal property
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\mathfrak{g}$$\scriptstyle{\rho_{U}}$$\scriptstyle{\psi_{f}}$$\scriptstyle{G(f)}$$\scriptstyle{\sigma}$
We will call this map $\Sigma_{f}\coloneqq\sigma\circ\psi_{f}$, which fulfills the cocycle condition, because $\psi_{f}$ is a group homomorphism and the properties of $\sigma$. Consider the diagram
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$G(\mathbb{C})$$\scriptstyle{\rho_{U}}$$\scriptstyle{\rho}$$\scriptstyle{G(f)}$$\scriptstyle{G(h_{\rho})}$
Let $f\in T_{\rho}G=\{f|R(\Gamma,G)\xrightarrow{f}k[\varepsilon,\eta]\xrightarrow{\pi}k,\pi\circ f=h_{\rho}\}$ then we get a morphism
$\Psi_{\rho}:\ T_{\rho}\mathrm{Hom}(\Gamma,G)$ $\longrightarrow Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$
$f$ $\longmapsto\Sigma_{f}$
Now let us also define an inverse. Let $\Sigma\in Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$
$\phi_{\Sigma}:\ \Gamma$ $\longrightarrow G(\mathbb{\varepsilon,\eta})$
$\gamma$ $\longmapsto\Sigma(\gamma)\cdot\rho(\gamma)$
We calculate:
$\phi_{\Sigma}(\gamma\delta)$ $=\Sigma(\gamma\delta)\cdot\rho(\gamma)$
$=\Sigma(\gamma)\cdot Ad_{\rho(\gamma)}\Sigma(\delta)\cdot\rho(\gamma\delta)$
$=\Sigma(\gamma)\rho(\gamma)\Sigma(\delta)\rho(\gamma)^{-1}\rho(\gamma)\rho(\delta)$
$=\Sigma(\gamma)\rho(\gamma)\Sigma(\delta)\rho(\delta)$
$=\phi_{\Sigma}(\gamma)\phi_{\Sigma}(\delta)$
So this is indeed a group homomorphism. By the universal property we get a morphism
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\scriptstyle{\rho_{U}}$$\scriptstyle{\phi_{\Sigma}}$$\scriptstyle{G(f_{\Sigma})}$
Recall that $\mathfrak{g}\cong ker(\pi)$ and $\rho(\gamma)\in G(\mathbb{C})$. It follows that $\pi\circ\phi_{\Sigma}=\rho$, because
$\pi(\phi_{\Sigma}(\gamma))=\pi(\Sigma(\gamma\cdot\rho(\gamma)))=\rho(\gamma).$
So, we get the map
$\Phi:\ Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$ $\longrightarrow T_{\rho}\mathrm{Hom}(\Gamma,G)$
$\Sigma$ $\longmapsto f_{\Sigma}$
The only thing remaining is that we have to prove, that $\Psi_{\rho},\Phi_{\rho}$ are inverse to each other. Applying $\Phi_{\rho}(\Psi_{\rho}(f))=f_{\Sigma_{f}}$. By the universal property
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\scriptstyle{\rho_{U}}$$\scriptstyle{\phi_{\Sigma_{f}}}$$\scriptstyle{G(f_{\Sigma_{f}})}$
we can now calculate
$\phi_{\Sigma_{f}}(\gamma)$ $=\Sigma_{f}(\gamma)\cdot\rho(\gamma)$
$=(\sigma\circ G(f)\circ\rho_{U})(\gamma)\cdot\rho(\gamma)$
$=(G(f)\circ\rho_{U}(\gamma))\cdot((\pi\circ G(f)\circ\rho_{U})(\gamma))^{-1}\cdot\pi\circ G(f)\circ\rho_{U})(\gamma)$
$\phi_{\Sigma_{f}}(\gamma)$ $=G(f)\circ\rho_{U}(\gamma)$
By uniqueness of the universal property, we get $\phi_{\Sigma_{f}}=\psi_{f}$. It follows that $G(f)=G(f_{\Sigma_{f}})$ so we get $f=f_{\Sigma_{f}}$. The other direction goes as follows:
$(\Psi_{\rho}\circ\Phi_{\rho})(\Sigma)$ $=\sigma\circ\psi_{f_{\Sigma}}$
$=\sigma\circ G(f_{\Sigma}\circ\rho_{U})$
$=(G(f_{\Sigma})\circ\rho_{U})(\gamma)((\pi\circ G(f_{\Sigma})\circ\rho_{U})(\gamma))^{-1}$
$=\phi_{\Sigma}(\gamma)\cdot\rho(\gamma)(\pi(\phi_{\Sigma}(\gamma)))^{-1}$
$=\Sigma(\gamma)\cdot\rho(\gamma)(\pi(\Sigma(\gamma)\cdot\rho(\gamma)))^{-1}$
$=\Sigma(\gamma)\cdot\rho(\gamma)\cdot e\cdot\rho(\gamma)^{-1}$
$=\Sigma(\gamma)$
So indeed the two morphisms are inverse to each other and we get the isomorphism
$Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})\cong T_{\rho}\mathrm{Hom}(\Gamma,G)$
The next proposition, was originally proven in [13] and we will follow the explicit exposition given in [19].
Proposition A.3.
Let $X$ be a Riemann surface of genus $g$ and $G\in\mathrm{Grp(Sch)}$, such that it admits a symmetric, bilinear, non-degenerate pairing. Then the dimension of the 1-cocycles is given by
(73) $\mathrm{dim}Z^{1}(\Gamma,Ad_{\rho})=(2g-1)\mathrm{dim}(G)+\mathrm{dim}(C_{G}(\pi_{1}(X))).$
Proof.
We give a sketch of the proof by providing the key arguments, which should convince the reader. First of all, since $X$ is an Eilenberg-MacLane space $K(\pi_{1}(X),1)$ computing group cohomology is the same as computing de Rham cohomology with local coefficients, which lets us identify
(74) $H^{*}(\pi_{1}(X),Ad_{\rho})\cong H^{*}_{dR}(X,E_{\rho}),$
so the cohomology vanishes in degrees larger than $2$. Furthermore, we have the Euler characteristic formula,
(75) $\chi(X,\rho)=\mathrm{dim}(H^{0}(\pi_{1}(X),Ad_{\rho}))-\mathrm{dim}(H^{1}(\pi_{1}(X),Ad_{\rho}))+\mathrm{dim}(H^{2}(\pi_{1}(X),Ad_{\rho})).$
This is independent of $\rho$, because when defining group cohomology, the twisting by $\rho$ is only realized inside of the differentials of the chain complex, so the chain complex itself does not see the twisting. For the trivial representation $\rho$ the Euler characteristic turns into
(76) $\chi(X,\mathrm{triv})=\chi(X)\mathrm{dim}(G).$
By Poincaré duality and the non-degeneracy of the Killing form, we obtain
(77) $H^{2}(\pi_{1}(X),Ad_{\rho})\overset{PD}{\cong}H^{0}(\pi_{1}(X),Ad_{\rho}^{\vee})^{\vee}\overset{Ad_{\rho}^{\vee}\cong Ad_{\rho}}{\cong}H^{0}(\pi_{1}(X),Ad_{\rho}).$
It is true, that $H^{0}(\pi_{1}(X),Ad_{\rho})$ is the space of $Ad(\rho)$-invariant elements and also that $\mathrm{dim}(B^{1}(\pi_{1}(X),Ad_{\rho}))=\mathrm{dim}(G)-\mathrm{dim}(C(\rho))$. Combining these two facts we obtain f
(78) $\begin{split}\mathrm{dim}(Z^{1}(\pi_{1}(X),Ad_{\rho}))&=\mathrm{dim}(H^{1}(\pi_{1}(X),Ad_{\rho}))+\mathrm{dim}(B^{1}(\pi_{1}(X),Ad_{\rho}))\\ &=\chi(X)\mathrm{dim}(G)+\mathrm{dim}(G)-\mathrm{dim}(C(\rho))\\ &=(2g-2)\mathrm{dim}(G)+\mathrm{dim}(G)-\mathrm{dim}(C(\rho))\\ &=(2g-1)\mathrm{dim}(G)+\mathrm{dim}(C(\rho))\end{split}$
We note that $C(\rho)=C(G)$ minimizes the dimension. ∎
As a corollary we obtain our wanted dimension of the tangent space.
Corollary A.4.
The dimension of the tangent space of a smooth point of $\mathrm{Rep}(\pi_{1}(X),\mathrm{OSp}(1|2))$ is $(2g-1)\mathrm{dim}(\mathrm{OSp}(1|2))$.
Proof.
A point is smooth if the dimension of the scheme and the Zariski tangent space coincide. Furthermore, the center of $\mathrm{OSp}(1|2)$ is given by $\{\pm\mathrm{id}\}$, thus the dimension of the center is 0 dimensional. By (A.3) the claim follows. ∎